OpenStudy (anonymous):

derivative of (x)/(x^2 -5) pls help me finish = (1 * (x^2 -5)- x *2x))/(x^2-5)^2

OpenStudy (anonymous):

how do i simplify further

OpenStudy (anonymous):

help anyone?

OpenStudy (anonymous):

the only thing I see is to sum the x^2 with the -2x^2, but the fraction I'd leave it like that, there is a way of dividing polynomials but I don't remember it now, and it doesn't always gives a perfect answer without rest.

OpenStudy (anonymous):

according to wolfram the final ans is -(x^2 + 5)/(x^2-5)^2 where did the -2x^2 go?

OpenStudy (anonymous):

Well, you need to sum with the x^2 as I said before, x^2-2x^2=-x^2

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

h-2h=-h, you have 1 x^2, and you take two x^2 out of it, then you get -1 x^2, it is simple subtraction